OpenStudy (deercult):

What is the height of an equilateral triangle with sides that are 12 cm long? Round to the nearest tenth.

OpenStudy (scoobydo):

6 because The height of an equilateral triangle forms a right triangle with one side as the hypotenuse and half of one side as a leg. The height is the other leg, so a is 1/2 length of side =1/2(12 cm)=6cm b is the height b=h c is the hypotenuse, one of the sides=12cm +a%5E2%2Bb%5E2=c%5E2, %286+cm%29%5E2%2Bh%5E2=%2812+cm%29%5E2 36+sq+cm+%2B+h%5E2=144+sq+cm subtract 36 sq cm from each side. h%5E2=108+cm Find square root of each side. h=10.4 cm ANSWER: The height is 10.4 cm. Answer B is correct. B.) s=length of side; d=diagonal d=sqrt%282%29s d=sqrt%282%29%28250+ft%29 d=1.414%28250+ft%29 d=353.6 feet ANSWER

OpenStudy (deercult):

Woah. Well, thanks for the answer. :0

OpenStudy (scoobydo):

your welcome

OpenStudy (deercult):

The quiz said it was 10.4 :"D But I got 6 working it out myself, so I'd have gotten it wrong anyways. Thanks again for trying!

OpenStudy (scoobydo):

so sorry

OpenStudy (deercult):

It's ok. At least I only missed two, and it was just a practice.